OpenStudy (anonymous):

Does y=12(1.3)^x represent exponential growth or exponential decay?

OpenStudy (lucaz):

as x increases the value of y also increases..

OpenStudy (anonymous):

I got it thanks anyway

OpenStudy (lucaz):

y=12(1.3)^x is the same as y=12(13/10)^x since the numerator 13 is greater than the denominator 10, this number increases as x increases